35. Harlan Corporation uses the periodic inventory system and the following information about their laptop computers is available:

Date Transaction Number of Units Cost per unit
01/01 Beginning Inventory 100 $800
05/05 Purchase 200 $900
08/10 Purchase 300 $1000
10/15 Purchase 200 $1050

During the year, 600 laptop computers were sold. What were ending inventory and cost of goods sold on 12/31 under the LIFO cost flow assumption?

(A) Ending inventory: $560,000 Cost of Goods Sold: $210,000

(B) Ending inventory: $210,000 Cost of Goods Sold: $560,000

(C) Ending inventory: $170,000 Cost of Goods Sold: $600,000

(D) Ending inventory: $600,000 Cost of Goods Sold: $170,000
Answer #1

Solution:

Ending Inventory units = (100 + 200+300+200) - 600 units = 200 units

Cost of ending inventory (LIFO method) = 100 units from Beginning + 100 units from 05/05 purchases

= 100 * $800 + 100 * $900 = $80,000 + $90,000 = $170,000

Cost of Goods sold =  200 units from 10/15 purchases +  300 units from 08/10 purchases + 100 units from 05/05 purchases

= 200 * $1050 + 300 *$1000 + 100*$900

= $210,000 + $300,000 + $90,000 = $600,000

Hence, option "C" is correct.